If I: = [0,4], calculate the norms of the following partitions:



a) P1: = (0,1,2,4)



b) P2: = (0,2,3,4)



c) P3: = (0,1,1.5,2,3.4,4)



d) P4: = (0,.5,2.5,3.5,4)

1
Expert's answer
2022-03-23T18:13:41-0400

Solution: The norm of a partition is merely the length of the largest subinterval into which the partition divides [a,b]. Clearly many partition have the same norm, so partition is not a function of the norm.

for "i=[a,b]" where "a<x_1<x_2<....<x_n<b," partition "p={x_1,x_2,....,x_n}" for above condition "norm=max\{x_2-x_1,x_3-x_2,...,x_n-x_{n-1}\}"


a)

"\therefore x_2-x_1=1-0=1
\\x_3-x_2=2-1=1
\\x_4-x_3=4-2=2
\\Hence ~ norm=max\{x_2-x_1,x_3-x_2, x_4-x_3\}
\\norm=max=\{1,1,2\}
\\norm=2"

Hence norm of the partition P1: = (0,1,2,4) is 2.


b)

"\therefore x_2-x_1=2-0=2
\\x_3-x_2=3-2=1
\\x_4-x_3=4-3=1
\\Hence ~ norm=max\{x_2-x_1,x_3-x_2, x_4-x_3\}
\\norm=max=\{2,1,1\}
\\norm=2"

Hence norm of the partition  P2: = (0,2,3,4) is 2.


c)

 "\therefore x_2-x_1=1-0=1
\\x_3-x_2=1.5-1=0.5
\\x_4-x_3=2-1.5=0.5
\\x_5-x_4=3.4-2=1.4
\\x_6-x_5=4-3.4=0.6
\\Hence ~ norm=max\{x_2-x_1,x_3-x_2, x_4-x_3,x_5-x_4,x_6-x_5\}
\\norm=max=\{1,0.5,0.5,1.4,0.6\}
\\norm=1.4"

Hence norm of the partition  P3: = (0,1,1.5,2,3.4,4) is 1.4.


d)

 "\therefore x_2-x_1=0.5-0=0.5
\\x_3-x_2=2.5-0.5=2
\\x_4-x_3=3.5-2.5=1
\\x_5-x_4=4-3.5=0.5
\\Hence ~ norm=max\{x_2-x_1,x_3-x_2, x_4-x_3,x_5-x_4\}
\\norm=max=\{0.5,2,1,0.5\}
\\norm=2"

Hence norm of the partition  P4: = (0,0.5,2.5,3.5,4) is 2.